• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Nilanga

    Log entry time 02:03:03 on July 6,2001

    Entry number 65635

    Followups:

    keyword=Left spectrometer helicity gate (Oh my god !)

    It seems that we have a serious problem with helicity gates on left spectroemter scalers (and on everything else on left spectrometer ?)

    According to the V-to-F scaler readings for left spectroemter every run seems to have a charge asymmetry of about 1500-2000 ppm ! ( 0.15%-0.2%), of course most of these runs had about 100 ppm level charge asymmetries and right spectroemter scalers seem to have this correct value. I checked many runs from 1455 to 1939 and all these runs seem to ahve the problem. I also checked different BCM signals (DX1, DX3, DX3 etc) and all the signals seem to be affected. So can this be a 1500 ppm level problem with the left spectrometer helicity gate?

    For runs with large charge asymmetry like run 1891/20891; with real asymmetries of about 7000 ppm, the two spectrometer scaler asymmetries still seem to differ by about 1500 ppm.

    At the begining of the experiment Bob checked charge asymmetry with the two spectrometer scalers and the BPM (using accelerator screen) and they all agreed; so this must have happened after that.